### v1

@@ Abstract
The best known fully retroactive priority queue costs $O(\log^{2}m\allowbreak\log\log m)$ time per operation, where $m$ is the number of operations performed on the data structure. In contrast, standard (non-retroactive) and partially retroactive priority queues cost $O(\log m)$ time per operation. So far, it is unknown whether this $O(\log m)$ bound can be achieved for fully retroactive priority queues.
In this work, we study a restricted variant of priority queues known as monotonic priority queues. We show that finding the minimum in a retroactive monotonic priority queue is a special case of the range-searching problem. We design a fully retroactive monotonic priority queue with a cost of $O(\log m+T(m))$ time per operation, where $T(m)$ is the maximum between the query and the update time of a specific range-searching data structure with $m$ elements. Finally, we design a fully retroactive monotonic priority queue that costs $O(\log m\log\log m)$ time per operation.
@@ 1 Introduction
  • A ‘standard’ bank system (data structure) can make new transactions (updates) in the present and show (query) the current balance.
  • A ‘retroactive’ bank system can modify past transactions (e.g., correct an old deposit), with all subsequent balances automatically updated to reflect this historical change.
  • A ‘partially retroactive’ bank system can make these historical changes, but still shows only the current balance.
  • Finally, a ‘fully retroactive’ bank system can both modify past transactions and show the account balance at any time (that is, after any transaction).
@@
Let $m$ be the number of operations performed in the data structure, Demaine et al. [6] designed a partially retroactive priority queue with $O(\log m)$ time per operation, matching the bounds of a binary heap [12]. In contrast, the most efficient fully retroactive priority queue so far has $O(\log^{2}m\log\log m)$ time per operation [7]. It is known that some fully retroactive data structures need to have a nonconstant multiplicative slowdown over its partially retroactive version [2], conditioned on well-believed conjectures. However, it is unknown whether this is the case for retroactive priority queues.
@@ 2.1 Definitions
  • Elements smaller than the last extracted cannot be inserted. [5].
@@
We define fully retroactive monotonic priority queues (FR-MonoPQ) in the same way as FR-PQ, but with the monotonic constraints applied.
@@ 2.2 Characteristics of Retroactive Priority Queues
(a) A retroactive priority queue with only $insert()$ operations.
@@
Using this representation, it becomes easy to see why a retroactive priority queue can be harder to implement efficiently than a standard one. Since once a retroactive update is made in the past, the extraction time of $O(m)$ elements can potentially change non-trivially. In order to propagate these changes, complex approaches are necessary, which can result in bad efficiency for the operations overall. Figure 2 shows an example of this. In this example, after the execution of the operation, a subset of elements had its extraction time changed. Therefore, all queries made after this insertion can potentially return a different value than what it would return before.
@@
Note that, in the partially retroactive version, since the past cannot be queried, we do not need to maintain the state of the priority queue at all times. Thus, there is more freedom in how we can propagate the changes caused by the operation. We only need to guarantee that the state of the priority queue is correct in the present. In the fully retroactive version, the ability to query at any point in time makes the process of updating the data structure while maintaining efficient queries more challenging, since it is apparently needed to maintain a correct priority queue state at all times.
@@ 2.3 Notation and Assumptions
For notational convenience, all elements and times are assumed to be real numbers. (However, the presented algorithms work for any totally ordered set that supports constant-time comparison, without affecting the asymptotic complexities.)
Assumption 2.6 (Unique Elements).
For simplicity, we assume that all elements inserted in the priority queue are distinct. (This assumption can be ignored if we do the following: Let $x$ be the element and $t$ be its insertion time. Instead of referring to $x$ directly, refer to it as $(x,t)$ instead. In this way, all elements become effectively unique again, since $t$ is unique (Assumption˜2.7).)
{observation}
Since all elements are unique, we have that $val[i]<val[j]$ for all $i<j$.
@@
We assume that if an element is extracted at time $t$, then the element does not exist at time $t$. In contrast, if an element is inserted at time $t$, then it is present in the priority queue at time $t$.
@@ 3.1 Conditions of Existence
An element $x$ exists only from its $\operatorname{insertionTime}(x)$ until just before its $\operatorname{extractionTime}(x)$. That is, $x$ only exists at times $t\in[\operatorname{insertionTime}(x),\operatorname{extractionTime}(x))$.
@@
Subsection˜3.1 is true because the only operation that modify the priority queue other than insertion is the $\operatorname{extract-min}()$ operation. Furthermore, since all inserted elements are distinct; the only way for them to stop existing is to be extracted; and they exist at time of insertion but not at time of extraction (Assumption˜2.9), it follows that Subsection˜3.1 is also true.
@@
Case 1: $x=\operatorname{lastExtracted}(t)$. Since $\operatorname{lastExtracted}(t)$ was extracted, it does not exist at time $t$. And $x$ exists at time $t$. Thus, they have the same value but are not the same element, contradicting that all elements are distinct (Assumption˜2.7).
@@ 4 The Data Structure
Finding $\operatorname{lastExtracted}(t)$ takes $O(\log m)$ time (Corollary˜3.6). And querying $T_{ins}$ takes $O(Q(m))$ time. Therefore, in total, $\operatorname{GetMin}(t)$ takes $O(\log m+Q(m))$ time.
@@ 5 Choosing a Range-Searching Data Structure
In the previous sections, we presented a FR-MonoPQ that uses a generic min-y range-searching data structure. We did that because we have not found a range-searching data structure for our specific use case of Definition˜3.10. In this section, we show that the 2D range tree of Mehlhorn and Näher [10, Thm. 7] can be used to answer queries in the form: Given an orthogonal rectangle, find the point inside this rectangle that has the smallest $y$-coordinate.
@@
6 Concluding Remarks
In this paper, we showed that finding the minimum in a retroactive monotonic priority queue is a special case of the range-searching problem. Then, we designed a FR-MonoPQ with $O(\log m)$ time per retroactive extraction, with $O(\log m+U(m))$ time per retroactive insertion, with $O(\log m+Q(m))$ time per retroactive query, and that uses $O(m\log m)$ space. Here $U(m)$, $Q(m)$ and $S(m)$ are the update, query, and space cost of a min-y range-searching data structure, respectively. Finally, we designed a FR-MonoPQ with $O(\log m)$ time per retroactive extraction, with $O(\log m\log\log m)$ time per retroactive query and insertion, and that uses $O(m\log m)$ space.

### v2

@@ Abstract
The best-known fully retroactive priority queue costs $O(\log^{2}m\allowbreak\log\log m)$ time per operation, where $m$ is the number of operations performed on the data structure. In contrast, standard (non-retroactive) and partially retroactive priority queues can cost $O(\log m)$ time per operation. So far, it is unknown whether this $O(\log m)$ bound can be achieved for fully retroactive priority queues.
In this work, we study a restricted variant of priority queues known as monotonic priority queues. First, we show that finding the minimum in a retroactive monotonic priority queue is a special case of the range-searching problem. Then, we design a fully retroactive monotonic priority queue with a cost of $O(\log m+T(m))$ time per operation, where $T(m)$ is the maximum between the query and the update time of a specific range-searching data structure with $m$ elements. Finally, we design a fully retroactive monotonic priority queue that costs $O(\log m\log\log m)$ time per operation.
@@
keywords:
computational geometry, monotone priority queue, range tree, retroactive data structure
@@ 1 Introduction
  • A standard bank system (data structure) can make new transactions (updates) in the present and show (query) the current balance.
  • A retroactive bank system can modify past transactions (e.g., correct an old deposit), with all subsequent balances automatically updated to reflect this historical change.
  • A partially retroactive bank system can make these historical changes but still shows only the current balance.
  • Finally, a fully retroactive bank system can both modify past transactions and show the account balance at any time (that is, after any transaction).
@@
Let $m$ be the number of operations performed in the data structure. Demaine et al. [6] designed a partially retroactive priority queue with $O(\log m)$ time per operation, matching the bounds of a binary heap [12]. In contrast, the most efficient fully retroactive priority queue so far has $O(\log^{2}m\log\log m)$ time per operation [7]. It is known that some fully retroactive data structures need to have a nonconstant multiplicative slowdown over its partially retroactive version [2], conditioned on well-believed conjectures. However, it is unknown whether this is the case for retroactive priority queues.
@@ 2.1 Definitions
  • Elements smaller than the last extracted cannot be inserted [5].
@@
We define fully retroactive monotonic priority queues (FR-MonoPQ) in the same way as FR-PQ but with the monotonic constraints applied.
@@ 2.2 Characteristics of Retroactive Priority Queues
(a) A retroactive priority queue with only insert operations.
@@
Using this representation, it becomes easy to see why a retroactive priority queue can be harder to implement efficiently than a standard one. Since once a retroactive update is made in the past, the extraction time of $O(m)$ elements can potentially change non-trivially. Thus, in order to propagate these changes, complex approaches are necessary, which can result in bad efficiency for the operations overall. Figure 2 shows an example of this. In this example, after the execution of the operation, a subset of elements had its extraction time changed. Therefore, all queries made after this insertion can potentially return a different value than what it would return before.
@@
Note that in the partially retroactive version, since the past cannot be queried, we do not need to maintain the state of the priority queue at all times. Thus, there is more freedom in how we can propagate the changes caused by the operation. We only need to guarantee that the state of the priority queue is correct in the present. In the fully retroactive version, the ability to query at any point in time makes the process of updating the data structure while maintaining efficient queries more challenging, since it is apparently needed to maintain a correct priority queue state at all times.
@@ 2.3 Notation and Assumptions
For notational convenience, all elements and times are assumed to be real numbers.
However, the presented algorithms work for any totally ordered set that supports constant-time comparison between two elements, without affecting the asymptotic complexities.
@@
Assumption 2.7 (Unique Elements).
For simplicity, we assume that all elements inserted in the priority queue are distinct.
This assumption can be ignored if we do the following: Let $x$ be the element and $t$ be its insertion time. Instead of referring to $x$ directly, refer to it as $(x,t)$ instead. In this way, all elements become effectively unique again, since $t$ is unique (Assumption˜2.6).
{observation}
Since all elements are unique, we have that $val[i]<val[j]$ for all $i<j$.
@@
We assume that if an element is extracted at time $t$, then the element does not exist at time $t$. Also, if an element is inserted at time $t$, then it is present in the priority queue at time $t$.
@@ 3.1 Conditions of Existence
An element $x$ exists only from its $\operatorname{insertionTime}(x)$ until just before its $\operatorname{extractionTime}(x)$. That is, $x$ only exists at times $t\in[\operatorname{insertionTime}(x),\allowbreak\operatorname{extractionTime}(x))$.
@@
Subsection˜3.1 is true because the only operation that modify the priority queue other than insertion is the $\operatorname{extract-min}()$ operation; and Subsection˜3.1 is true because the only way for the elements to stop existing is to be extracted (Subsection˜3.1), they exist at time of insertion but not at time of extraction (Assumption˜2.9), and all elements are distinct (Assumption˜2.7).
@@
Case 1: $x=\operatorname{lastExtracted}(t)$. Since $\operatorname{lastExtracted}(t)$ was extracted, it does not exist at time $t$. And $x$ exists at time $t$. Thus, they have the same value but are not the same element, contradicting that all elements are distinct (Assumption˜2.6).
@@ 4 The Data Structure
Finding $\operatorname{lastExtracted}(t)$ takes $O(\log m)$ time (Lemma˜3.6). And querying $T_{ins}$ takes $O(Q(m))$ time. Therefore, in total, $\operatorname{GetMin}(t)$ takes $O(\log m+Q(m))$ time.
@@ 5 Choosing a Range-Searching Data Structure
In this section, for the purpose of satisfying Definition˜3.10, we show that the 2D range tree of Mehlhorn and Näher [10, Thm. 7] can be used to answer queries in the form: Given an orthogonal rectangle, find the point inside this rectangle that has the smallest $y$-coordinate. Later, we directly use this range tree in Theorem˜4.1 to obtain new bounds for the FR-MonoPQ.
@@
6 Conclusion
In this paper, we showed that finding the minimum in a retroactive monotonic priority queue is a special case of the range-searching problem. Then, we designed a FR-MonoPQ with $O(\log m)$ time per retroactive extraction, with $O(\log m+U(m))$ time per retroactive insertion, with $O(\log m+Q(m))$ time per retroactive query, and that uses $O(m+S(m))$ space. Finally, we designed a FR-MonoPQ with $O(\log m)$ time per retroactive extraction, with $O(\log m\log\log m)$ time per retroactive query and insertion, and that uses $O(m\log m)$ space.